OpenStudy (anonymous):

solve this mental question: 15022=2+3+5 183652=4+2+9 482466=3+6+8 202541=5+4+5 then 5+2+7= ?

OpenStudy (amistre64):

define "mental"

OpenStudy (amistre64):

\[\color{white}{this is white}\]

OpenStudy (saifoo.khan):

it will start from 14.

OpenStudy (saifoo.khan):

14-35

OpenStudy (saifoo.khan):

Dont know the last two!

OpenStudy (anonymous):

7 + 2 + 5 = 143568 a + b + c = ab followed by ac followed by a(b+c) - b

OpenStudy (turingtest):

What's the pattern you see meverett? I see that this one is (ab)(ac)... but a(b+c)-b=7(2+5)-2=47 how does that work?

OpenStudy (anonymous):

opps! Thanks TT